OpenStudy (anonymous):

Suppose you have a set of data. The min is 300, the max is 1,700. Suppose then you add 150 to every single data point. The range is currently 1,400. By adding 150 to every single data point, how will this affect the range?

OpenStudy (anonymous):

@Mertsj @Zarkon @dpaInc @jim_thompson5910 @lgbasallote @KingGeorge @AccessDenied @SmoothMath @eliassaab @asnaseer @.Sam. @kropot72 @eigenschmeigen @satellite73 @amistre64 @TuringTest @shivam_bhalla @experimentX @FoolForMath @Callisto @phi

OpenStudy (anonymous):

So then what is the range?

OpenStudy (anonymous):

The range would be the same as it were before right? I know that that is true, but I am trying to show that

OpenStudy (anonymous):

Never mind, i misunderstood the question. Nothing happens.

OpenStudy (anonymous):

Consider: min = 3 max = 5 range = 2 You add +2 to each point, then min and max become: min = 5 max = 7 Then, you get the range, which equals 2 ------------------------------------------- But the situation I have is this: max = 1700 min = 300 range = 1400 Then, we add +150 to every data point, then min and max become: max = 1850 min = 450 range = 1400 Wow! Something is wrong with my calculator, or I am not typing it in properly.

OpenStudy (anonymous):

You are right, its the same as before look at this: \[Min:300\qquad Mx:1700\] the range is their difference. Look at the new data set: \[Min:450\qquad Mx:1850\] if you take the difference of your new data bounds it will be exactly the same as the one from before.

OpenStudy (mertsj):

It won't affect the range

OpenStudy (callisto):

If you add a constant (k) to every datum in a set of data , the range doesn't change... Originally, Min: x1 Max: x2 After addition Min: x1 +k Max: x2 +k Range = (x2 +k) - (x1+k) = x2 - x1 = same as the original range...

OpenStudy (kropot72):

The new minimum will be 300 + 150 = 450 The new maximum will be 1700 + 150 = 1850 The new range will be 1850 - 450 = 1400

OpenStudy (anonymous):

Thank you all kindly for your replies.

OpenStudy (anonymous):

I realized I made a calculation mistake.